Grade 11General Physics

the magnitudes of two vectors A and B are A = 12 units and B = 4 units calculate1 their resultant if the angle between them is 60 angle and 120 angle

matiwos
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student
 
As we know
 
R² = P² + Q² + 2PQ cos @
 
Hence
i) R² = 12² + 4² + 2*12*4 * cos60°
 
 R² = 144 + 16 + 96 * 1/2
 
 R² = 160 + 48 
 R² = 208
 
R = sqrt(208)
 
Now you can do calculation.
 
ii)
R² = 144 + 16 + 96 cos 120°
R² = 160 - 96 *1/2
R² = 160-48
 
R² = 112
 
R = sqrt(112)
 
 
Regards
Arun (askIITians forum expert)